OpenStudy (anonymous):

dy/dx=cos^2(x)cos^2(2y), y=?

OpenStudy (anonymous):

This is separable. You just need to rewrite it as\[\frac{dy}{\cos^2 (2y)}=\cos^2 (2x) dx \rightarrow \int\limits_{}{} \frac{dy} {\cos^2 (2y)}=\int\limits_{}{} \cos^2 (2x) dx\]

OpenStudy (anonymous):

Use the fact that \[\cos 2\theta =\cos^2 \theta - \sin^2 \theta = 2 \cos^2 \theta -1 \rightarrow \cos^2\theta =\frac{\cos 2 \theta +1}{2}\]

OpenStudy (anonymous):

Hello dichalao

OpenStudy (anonymous):

Hello loki~ haha

OpenStudy (anonymous):

Isn't it ~1am there?

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

i dont want to end my day as usual :(

OpenStudy (anonymous):

Why's that?

OpenStudy (anonymous):

find the exact solution to this bvp u''-u=0 u(0)=0 and u(1)=1 how do you solve this problem..i was trying to help the guy, but failed

OpenStudy (anonymous):

I don know.. i just dont want to go to bed

OpenStudy (anonymous):

It should be a exponential function right

OpenStudy (anonymous):

Lol...this is a linear homogeneous second order differential equation with constant coefficients. We assume a solution of the form \[y=e^{\lambda x}\], sub it in and solve the characteristic equation\[{\lambda}^2-1=0 \rightarrow \lambda = \pm 1\]The solution's then\[y=c_1 e^x + c_2 e^{-x}\]

OpenStudy (anonymous):

how you got λ2−1=0→λ=±1

OpenStudy (anonymous):

\[(e^{\lambda x})''-e^{\lambda x}=0 \rightarrow {\lambda}^2e^{\lambda x}-e^{\lambda x}=0 \rightarrow ({\lambda}^2-1)e^{\lambda x}=0\]In the last product, it can only be the case that\[{\lambda}^2-1=0\]since\[e^{\lambda x} \neq0 \]for all x.

OpenStudy (anonymous):

because λ=+-1 you got ce^x+c2e^(-x)?

OpenStudy (anonymous):

Yep, the aim is to find lambdas that will allow you to make the assumption true. There's a theorem in differential equations that tells us how many solutions there will be in a diff. equation and that those solutions are unique. For second order homogeneous (what we have here) there will be two solutions, and since we've found two independent solutions (that technically needs to be checked with something called the Wronskian, but no-one usually cares with these types), and since we know a set of solutions is unique, we've found the solutions to the equation.

OpenStudy (anonymous):

Last bit wasn't explained all that well.

OpenStudy (anonymous):

LOL it is good enough at my level

OpenStudy (anonymous):

Happy now? You can hit the sheets!

OpenStudy (anonymous):

LOL i am always happy when you are on :) flattery gets me to anywehre

OpenStudy (anonymous):

So do you just have to tell that guy the solution?

OpenStudy (anonymous):

you can do it :) i dont want to take any credits for it.. it is 10 thread up

OpenStudy (anonymous):

You can just screen shot it and send as an attachment.

OpenStudy (anonymous):

He just has to solve for the boundary conditions u(0)=0 and u(1)=1.

OpenStudy (anonymous):

LOL can you please do that for me :P

OpenStudy (anonymous):

so you will earn your 133th fan

OpenStudy (anonymous):

Okay, I'll go over to him. I think nikvist might be dealing with it...